Hello, I have tried to figure this out but with no success can you help.
Bob invested $20,000, part at 14% and part at 13%. If the total interest at the end of the year is $2,720, how much did he invest at 14%?
this is what I have done so far but it is not proofing out. What am I doing wrong
0.14x+0.13(20,000-x)=2,720
0.14x+2,800-.013x=2,720
0.13x=2,720-2,800
.013x=200
X=200/0.13
X=1538.46 invested @ 14%
20,000-923.07 = 19,076.93 invested @ 13%

Answer by rapaljer(4671)

You seem to have TWO errors in your second step!
0.14x+0.13(20,000-x)=2,720
0.14x+2,800-.013x=2,720

Should be
0.14x+0.13(20,000-x)=2,720
0.14x+2,600-0.13x=2,720

Continuing:
0.01x = 120
x= $12000
20000-x = $8000

Check: .14(12000)+.13(8000)=2720

R^2